# Mixture-of-RAG: Integrating Text and Tables with Large Language Models

### v2

@@ Abstract
Large language models (LLMs) achieve optimal utility when their responses are grounded in external knowledge sources. However, real-world documents, such as annual reports, scientific papers, and clinical guidelines, frequently combine extensive narrative content with complex, hierarchically structured tables. While existing retrieval-augmented generation (RAG) systems effectively integrate LLMs’ generative capabilities with external retrieval-based information, their performance significantly deteriorates especially processing such heterogeneous text-table hierarchies. To address this limitation, we formalize the task of Heterogeneous Document RAG, which requires joint retrieval and reasoning across textual and hierarchical tabular data. We propose MixRAG, a novel three-stage framework: (i) hierarchy row-and-column-level (H-RCL) representation that preserves hierarchical structure and heterogeneous relationship, (ii) an ensemble retriever with LLM-based reranking for evidence alignment, and (iii) multi-step reasoning decomposition via a RECAP prompt strategy. To bridge the gap in available data for this domain, we release a large-scale dataset DocRAGLib, a 2k-document corpus paired with automatically aligned text-table summaries and gold document annotations. The comprehensive experiment results demonstrate that MixRAG boosts top-1 retrieval by 46% over strong text-only, table-only, and naive-mixture baselines, establishing new state-of-the-art performance for mixed-modality document grounding.
1 Introduction
@@
With the rapid development of large language models (LLMs) [20; 14], the reliability of their response substantially depends on the extent to which they can be grounded in verifiable evidence. Retrieval-Augmented Generation (RAG) has emerged as an effective strategy that combines the generative power of LLMs with external retrieval-based information. Recent research has extended the application of RAG to various scenarios, including knowledge graph [24; 17], database [2], and multi-modal data [15; 32]. When the external source is plain text or simple tables, RAG systems retain precision and fluency. However, real-world data, including financial statements, clinical guidelines, research articles, and governmental reports, all interleave complex hierarchical tables whose rows and columns nest multiple semantic layers. The crucial metrics numbers, such as earnings per share, dosage limits, and experimental results, are embedded within these tables, and the relationships among those numbers are defined by the headers that span multiple rows or columns. Therefore, the research on the Heterogeneous Document RAG remains nascent. The Heterogeneous Document RAG aims to efficiently retrieve and synthesize relevant information from large, complex documents by integrating textual and hierarchical tabular data with the power of LLMs for more comprehensive generation.
Figure 1: An Example of Heterogeneous Document RAG Task
@@
The Heterogeneous Document RAG task involves retrieving the most relevant documents from extensive document repositories based on a specific question and generating the corresponding answer. However, it is challenging due to the complexity and scale of real-world repositories, which often contain numerous indistinguishable documents with intricate structures. For example, Statistics Canada offers thousands of statistical reports across subjects, including 1,761 on "Labour" and 1,132 on "Population and demography", with semantically similar content often combining textual descriptions and hierarchical tables.
@@
Some novel RAG techniques  [13; 3; 21; 12] have been explored to improve retrieval accuracy by enhancing RAG frameworks. While effective in text-based tasks, these methods often fail to capture the critical table information within retrieved documents. Other table-focused RAG techniques [2; 18; 10; 17] lack seamless integration of table data with text, resulting in unaligned representations and difficulty handling complex hierarchical structures. As shown in Figure 1, when retrieval relies on the hierarchical structure of a table, such methods often break down the complex structure, leading to information loss, i.e., disruption of critical dependencies (e.g., 2012-2014 trends), and a reduced ability to capture critical relationships within the data. Additionally, while well-established document-based QA methods [27; 19; 26] have demonstrated relatively high accuracy, they still struggle with complex tasks involving documents with both long text and hierarchical tables [31]. (More related work attached in Appendix I.) Overall, although existing methods have provided preliminary solutions for Heterogeneous Document RAG, achieving more precise information retrieval and integration still faces numerous challenges: (i) Data Scarcity. The publicly available benchmarks lack a cleaned, integrated corpus pairing narrative text with hierarchical tables. (ii) Structured Representation. The unified representation must preserve row-column hierarchies while remaining consistent with text. As shown in Figure 1, the key terms "Total Buy" and "Revenues of Electric Sales" are strongly linked to the hierarchical structure within the heterogeneous document, and the information contained within tables is dense, highlighting the challenge of representing structured data in such documents. (iii) Retrieval Precision. The traditional semantic similarity-based retrieval methods often return text chunks that, while semantically related to the question, may not provide valid evidence for reasoning. For example, as illustrated in Figure 1, existing methods may retrieve documents containing the exact year ’2014’ from the query, while missing the relevant historical data from 2012 and 2013 that the query actually seeks. (iv) Reasoning Accuracy. Effective retrieval requires hierarchical alignment between unstructured text and structured tabular data, necessitating the ability to process and reason across heterogeneous documents accurately (e.g., deriving 2012-2013 context from a 2014 query and multi-step computation).
@@
To address the above challenges in Heterogeneous Document RAG, we propose the DocRAGLib dataset and MixRAG framework. Specifically, DocRAGLib consists of a document repository with 2,178 documents, each of which combines text and hierarchical tables, as well as 4,468 QA pairs. DocRAGLib is collected from reliable sources and cleaned to ensure consistency across document types and formats. To the best of our knowledge, DocRAGLib is the first dataset that bridges the gap of heterogeneous documents. MixRAG framework consists of three modules: Heterogeneous Document Representation Module, Cross-Modal Retrieval Module, and Multi-Step Reasoning Module. Specifically, the Heterogeneous Document Representation Module employs a hierarchical row-and-column-level (H-RCL) table summarization method to capture the structure and content of complex tables. It generates representations that preserve table structures and are optimized for retrieval, addressing the challenge of effectively representing table information in heterogeneous documents. The Cross-Modal Retrieval Module overcomes the limitations of relying solely on semantic similarity retrieval by adopting a two-stage approach. First, the ensemble retrieval stage combines BM25 with the semantic understanding of embedding retrieval to filter candidate documents. Second, the LLM-based reranking stage utilizes the contextual reasoning ability of LLMs to identify the most relevant document, improving the accuracy and comprehensiveness of retrieval. The Multi-Step Reasoning Module introduces the RECAP prompt strategy, which decomposes complex reasoning into sub-tasks and leverages external calculators to manage mathematical operations during the reasoning process. MixRAG significantly improves answer accuracy and effectively addresses the challenges of multi-step reasoning and complex calculations in heterogeneous documents.
@@
  • We propose a new large-scale dataset, DocRAGLib, which provides a high-quality benchmark for heterogeneous information retrieval tasks.
@@
2 Preliminary
@@ 2.1 Task Definition
Document Corpus. A comprehensive collection of documents $\mathbf{C}=\{D_{1},D_{2},\dots,D_{|\mathbf{C}|}\}$ where each document $D_{i}\in\mathbf{C}$ consists of multiple text segments $\mathbf{P}=\{P_{1},P_{2},\dots,P_{|\mathbf{P}|}\}$ and several complex hierarchical tables $\mathbf{T}=\{T_{1},T_{2},\dots,T_{|\mathbf{T}|}\}$, referred to as heterogeneous document. Each hierarchical table is defined as a tuple $T=\{{H}_{t},{H}_{l},d\}$, where ${H}_{t}$ denotes the top headers in hierarchical structure of table $T$, ${H}_{l}$ denotes the left headers in hierarchical structure of table $T$ (example of hierarchical headers shown in Figure 1), and $d$ denotes the data cells of table $T$.
Question. A question $Q$ posed by the user, for which there exists a unique document of highest relevance $D^{*}\in\mathbf{C}$ in the corpus.
@@
$A=\mathcal{F}(Q,\mathbf{C})=\text{Inference}\left(Q,D^{*}\right),\text{where }D^{*}=\text{Retrieve}(Q,\mathbf{C})$ (1)
@@
2.2 Dataset
To bridge the gap of heterogeneous data availability, we construct the DocRAGLib dataset based on existing relevant datasets  [30; 9] and data crawled from the web. DocRAGLib contains 2,178 documents, combining text and table content, and includes 4,468 QA pairs designed based on the content of these documents. Existing datasets are predominantly tailored for single-document RAG tasks, limiting their applicability to scenarios requiring the processing of large-scale and complex document corpora. DocRAGLib aims to address this limitation by providing a comprehensive dataset specifically designed for multi-document RAG tasks. The details of Data Collection, Data Cleaning, and Data Statistics are elaborated in Appendix A.
3 Methodology
@@
Figure 2: The Overview of our MixRAG Framework
@@
3.1 Heterogeneous Document Representation Module
The Heterogeneous Document Representation Module focuses on generating effective representations of heterogeneous documents, capturing both lengthy text and complex hierarchical tables to optimize retrieval performance. A traditional approach, the Table Level Summary, involves providing the table’s schema to the LLM and summarizing it. However, it oversimplifies the table, leading to the loss of key information and ignoring important structural details. For effective representation, the Heterogeneous Document Representation Module first introduces a row-and-column-level(RCL) table summary to represent tables. Then, leveraging the hierarchical structure of the tables, we extend this RCL table summary to capture the nested relationships and dependencies within the table data, ensuring a more comprehensive representation (see Appendix B for examples and comparison of different representation strategies).
@@ 3.1 Heterogeneous Document Representation Module
$\left\{\begin{aligned} H_{t}&=\{H_{t}^{r}\mid r=1,\dots,R_{t}\}\\ H_{l}&=\{H_{l}^{r}\mid r=1,\dots,R_{l}\}\end{aligned}\right$. (2)
, where $R_{t}$ denotes the number of hierarchical levels in the top headers, and $R_{l}$ denotes the number of hierarchical levels in the left headers.
@@
, where $h_{t}^{r}(j)$ represents the header at level $r$ in the top header hierarchy, while $h_{l}^{r}(i)$ represents the header at level $r$ in the left header hierarchy. As shown in Figure 3, these headers form multi-level paths used to locate specific data cells within the table. To pinpoint a particular data cell $d_{ij}$, we define the paths from both the left and top headers, $P_{l}(i)$ and $P_{t}(j)$, respectively, which together uniquely identify the location of the data cell:
$\left\{\begin{aligned} P_{l}(i)&=h_{l}^{1}(i_{1})\to h_{l}^{2}(i_{2})\to\dots\to h_{l}^{R_{l}}(i_{R_{l}})\\ P_{t}(j)&=h_{t}^{1}(j_{1})\to h_{t}^{2}(j_{2})\to\dots\to h_{t}^{R_{t}}(j_{R_{t}})\end{aligned}\right$. (4)
@@
3.1.1 General RCL Table Summary

### v3

@@ Abstract.
Large language models (LLMs) achieve optimal utility when their responses are grounded in external knowledge sources. However, real-world documents, such as annual reports, scientific papers, and clinical guidelines, frequently combine extensive narrative content with complex, hierarchically structured tables. While existing retrieval‑augmented generation (RAG) systems effectively integrate LLMs’ generative capabilities with external retrieval-based information, their performance significantly deteriorates especially processing such heterogeneous text-table hierarchies. To address this limitation, we formalize the task of Heterogeneous Document RAG, which requires joint retrieval and reasoning across textual and hierarchical tabular data. We propose MixRAG, a novel three‑stage framework: (i) hierarchy row-and-column-level (H-RCL) representation that preserves hierarchical structure and heterogeneous relationship, (ii) an ensemble retriever with LLM-based reranking for evidence alignment, and (iii) multi‑step reasoning decomposition via a RECAP prompt strategy. To bridge the gap in available data for this domain, we release the dataset DocRAGLib, a 2k‑document corpus paired with automatically aligned text‑table summaries and gold document annotations. The comprehensive experiment results demonstrate that MixRAG boosts top‑1 retrieval by 46% over strong text‑only, table‑only, and naive‑mixture baselines, establishing new state-of-the-art performance for mixed-modality document grounding.
Information Retrieval, Question Answering, Retrieval-Argument Generation
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 9–13, 2026; Jeju Island, Republic of Korea.booktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD 2026), August 9–13, 2026, Jeju Island, Republic of Koreaisbn: 979-8-4007-2258-5/2026/08doi: 10.1145/3770854.3780171ccs: Information systems Information retrievalccs: Computing methodologies Natural language generation
1. Introduction
@@
With the rapid development of large language models (LLMs) (llama2; mistral7b), the reliability of their response substantially depends on the extent to which they can be grounded in verifiable evidence. Retrieval-Augmented Generation (RAG) has emerged as an effective strategy that combines the generative power of LLMs with external retrieval-based information. Recent research has extended the application of RAG to various scenarios, including knowledge graph (Graph_MDQA; matsumoto2024kragen), database (biswal2024tag), and multi-modal data (multi-moda_PreFLMR; Multi-moda-sigir1). When the external source is plain text or simple tables, RAG systems retain precision and fluency. However, real-world data, including financial statements, clinical guidelines, research articles, and governmental reports, all interleave complex hierarchical tables whose rows and columns nest multiple semantic layers. The crucial metrics numbers, such as earnings per share, dosage limits, and experimental results, are embedded within these tables, and the relationships among those numbers are defined by the headers that span multiple rows or columns. Therefore, the research on the Heterogeneous Document RAG remains nascent. The Heterogeneous Document RAG aims to efficiently retrieve and synthesize relevant information from large, complex documents by integrating textual and hierarchical tabular data with the power of LLMs for more comprehensive generation.
Figure 1. An Example of Heterogeneous Document RAG Task
@@
The Heterogeneous Document RAG task involves retrieving the most relevant documents from extensive document repositories based on a specific question and generating the corresponding answer. However, it is challenging due to the complexity and scale of real-world repositories, which often contain numerous indistinguishable documents with intricate structures. For example, Statistics Canada offers thousands of statistical reports across subjects, including 1,761 on “Labour” and 1,132 on Population and “demography”, with semantically similar content often combining textual descriptions and hierarchical tables.
@@
Some novel RAG techniques  (izacard2020distilling; borgeaud2022improving; wang-etal-2024-dapr; islam-etal-2024-open) have been explored to improve retrieval accuracy by enhancing RAG frameworks. While effective in text-based tasks, these methods often fail to capture the critical table information within retrieved documents. Other table-focused RAG techniques (biswal2024tag; roychowdhury2024eratta; cheng2024retrievalonmultimodal; matsumoto2024kragen) lack seamless integration of table data with text, resulting in unaligned representations and difficulty handling complex hierarchical structures. As shown in Figure 1, when retrieval relies on the hierarchical structure of a table, such methods often break down the complex structure, leading to information loss, i.e., disruption of critical dependencies (e.g., 2012-2014 trends), and a reduced ability to capture critical relationships within the data. Additionally, while well-established document-based QA methods (ProTrix; srivastava-etal-2024-evaluating; TACR) have demonstrated relatively high accuracy, they still struggle with complex tasks involving documents with both long text and hierarchical tables (zhao-etal-2024-docmath). Overall, although existing methods have provided preliminary solutions for Heterogeneous Document RAG, achieving more precise information retrieval and integration still faces numerous challenges: (i) Data Scarcity. The publicly available benchmarks lack a cleaned, integrated corpus pairing narrative text with hierarchical tables. (ii) Structured Representation. The unified representation must preserve row‑column hierarchies while remaining consistent with text. As shown in Figure 1, the key terms “Total Buy” and “Revenues of Electric Sales” are strongly linked to the hierarchical structure within the heterogeneous document, and the information contained within tables is dense, highlighting the challenge of representing structured data in such documents. (iii) Retrieval Precision. The traditional semantic similarity-based retrieval methods often return text chunks that, while semantically related to the question, may not provide valid evidence for reasoning. For example, as illustrated in Figure 1, existing methods may retrieve documents containing the exact year ‘2014’ from the query, while missing the relevant historical data from 2012 and 2013 that the query actually seeks. (iv) Reasoning Accuracy. Effective retrieval requires hierarchical alignment between unstructured text and structured tabular data, necessitating the ability to process and reason across heterogeneous documents accurately (e.g., deriving 2012-2013 context from a 2014 query and multi-step computation).
@@
To address the above challenges in Heterogeneous Document RAG, we propose the DocRAGLib dataset and MixRAG [^1] Our dataset, code, and prompts are available in GitHub at https://github.com/ChiZhang-bit/Mixture-of-RAG. framework. Specifically, DocRAGLib consists of a document repository with 2,178 documents, each of which combines text and hierarchical tables, as well as 4,468 QA pairs. DocRAGLib is collected from reliable sources and cleaned to ensure consistency across document types and formats. To the best of our knowledge, DocRAGLib is the first dataset that bridges the gap of heterogeneous documents. MixRAG framework consists of three modules: Heterogeneous Document Representation Module, Cross-Modal Retrieval Module, and Multi-Step Reasoning Module. Specifically, the Heterogeneous Document Representation Module employs a hierarchical row-and-column-level (H-RCL) table summarization method to capture the structure and content of complex tables. It generates representations that preserve table structures and are optimized for retrieval, addressing the challenge of effectively representing table information in heterogeneous documents. The Cross-Modal Retrieval Module overcomes the limitations of relying solely on semantic similarity retrieval by adopting a two-stage approach. First, the ensemble retrieval stage combines BM25 with the semantic understanding of embedding retrieval to filter candidate documents. Second, the LLM-based reranking stage utilizes the contextual reasoning ability of LLMs to identify the most relevant document, improving the accuracy and comprehensiveness of retrieval. The Multi-Step Reasoning Module introduces the RECAP prompt strategy, which decomposes complex reasoning into sub-tasks and leverages external calculators to manage mathematical operations during the reasoning process. MixRAG significantly improves answer accuracy and effectively addresses the challenges of multi-step reasoning and complex calculations in heterogeneous documents.
@@
  • We propose a new dataset, DocRAGLib, which provides a high-quality benchmark for heterogeneous information retrieval tasks.
@@
2. Related Work
In this section, we present an overview of relevant research from two perspectives: RAG and Document QA.
2.1. RAG
RAG has significantly enhanced language models, particularly in knowledge-intensive tasks involving document data. This section presents existing work from two main aspects: advancements in RAG techniques (borgeaud2022improving; wang-etal-2023-query2doc; wang-etal-2024-dapr; izacard2020distilling; yan2024crag; islam-etal-2024-open; asai2023selfrag; cheng-etal-2024-call) and its applications across different modalities (biswal2024tag; cheng2024retrievalonmultimodal; matsumoto2024kragen; roychowdhury2024eratta; hu2024mrag).
Advancements in RAG techniques aim to improve retrieval accuracy, efficiency, and robustness in knowledge-intensive tasks. Izacard et al. (izacard2020distilling) use knowledge distillation to train retrievers without annotated query-document pairs, relying on reader model attention for synthetic labels, though it depends on the reader model’s quality. RETRO (borgeaud2022improving) combines document chunks retrieved by local similarity with preceding tokens in auto-regressive models but faces challenges with long documents. DAPR (wang-etal-2024-dapr) addresses this by integrating hybrid retrieval with BM25 and contextualized passage representations for long document passages. Open-RAG (islam-etal-2024-open) improves multi-hop reasoning and retrieval accuracy by transforming dense LLMs into the parameter-efficient sparse mixture of expert models. However, these methods primarily focus on text-based RAG, with limited progress in handling mixed modalities, which pose unique retrieval and reasoning challenges.
Recent advancements in RAG have extended to multimodal data beyond text. TAG (biswal2024tag) and ERATTA (roychowdhury2024eratta) introduce RAG frameworks for table-structured data. Chen et al. (chen-etal-2024-table) propose a method combining join relationship discovery and mixed-integer programming-based re-ranking for table retrieval. RAGTrans (cheng2024retrievalonmultimodal) integrates textual and multimedia information for enhanced representation and retrieval. KRAGEN (matsumoto2024kragen) introduces a knowledge graph-enhanced RAG framework that uses graph-based retrieval to improve factual consistency and reasoning in biomedical problems. MRAG-Bench (hu2024mrag) focuses on vision-centric RAG, highlighting cases where visual knowledge outperforms textual information, especially for vision-language models. However, none of these approaches address RAG for the combined modality of tables and text within documents, a gap that MixRAG specifically addresses.
2.2. Document QA
Document QA refers to the task of extracting or generating answers based on the content of a given document. This section reviews existing approaches to Document QA from two primary perspectives: methods leveraging pre-trained language models (chen-etal-2020-hybridqa; zhu-etal-2021-tat; cheng-etal-2022-hitab; zhao-etal-2022-multihiertt; 10.1007/978-3-031-44693-1_46) and methods utilizing LLM (luo2023hrot; srivastava-etal-2024-evaluating; CoT; PoT; zhao-etal-2024-docmath).
In Document QA, pre-trained NLP models are fine-tuned to process document-question pairs, enabling them to extract information or generate answers based on the content. HybridQA (chen-etal-2020-hybridqa) and TAT-QA (zhu-etal-2021-tat) apply QA tasks to text-table hybrid documents, but their tables are typically flat. HiTab (cheng-etal-2022-hitab) focuses on hierarchical table-based QA tasks but struggles with hybrid data combining text and hierarchical tables. Zhao et al. (zhao-etal-2022-multihiertt) introduce the MultiHiertt benchmark, containing both text and hierarchical tables and propose MT2Net for this task. MT2Net extracts facts from hybrid documents and performs reasoning, but it has high content requirements and depends heavily on table cell descriptions for accurate reasoning. NAPG (10.1007/978-3-031-44693-1_46) outperforms MT2Net on MultiHiertt, yet existing NLP models still face challenges with multi-step reasoning, especially in complex mathematical computations.
With the emergence of advanced LLMs, several studies apply them to question-answering tasks with hybrid documents. Luo et al. (luo2023hrot) introduce the HRoT strategy for text-table hybrid QA, achieving better results than existing pre-trained NLP models on the MultiHiertt dataset. Srivastava et al. (srivastava-etal-2024-evaluating) propose EEDP, designed for semi-structured documents, and compare it with other prompting methods (such as CoT (CoT) and PoT (PoT)) across various QA datasets. Zhao et al. (zhao-etal-2024-docmath) evaluate the numerical reasoning abilities of 27 LLMs using CoT and PoT on text-table hybrid documents. Despite the progress of advanced prompting strategies in guiding LLMs for hybrid document QA, there remains room for improvement compared to human performance, highlighting the need for more efficient prompting techniques.
3. Preliminary
@@ 3.1. Task Definition
1) Document Corpus. A comprehensive collection of documents $\mathbf{C}=\{D_{1},D_{2},\dots,D_{|\mathbf{C}|}\}$ where each document $D_{i}\in\mathbf{C}$ consists of multiple text segments $\mathbf{P}=\{P_{1},P_{2},\dots,P_{|\mathbf{P}|}\}$ and several complex hierarchical tables $\mathbf{T}=\{T_{1},T_{2},\dots,T_{|\mathbf{T}|}\}$, referred to as heterogeneous document. Each hierarchical table is defined as a tuple $T=\{{H}_{t},{H}_{l},d\}$, where ${H}_{t}$ denotes the top headers in hierarchical structure of table $T$, ${H}_{l}$ denotes the left headers in hierarchical structure of table $T$ (example of hierarchical headers shown in Figure 1), and $d$ denotes the data cells of table $T$.
2) Question. A question $Q$ posed by the user, for which there exists a unique document of highest relevance $D^{*}\in\mathbf{C}$ in the corpus.
@@
(1) $\small A=\mathcal{F}(Q,\mathbf{C})=\text{Inference}\left(Q,D^{*}\right),\text{where }D^{*}=\text{Retrieve}(Q,\mathbf{C})$
@@
3.2. Dataset
Existing datasets are predominantly tailored for single-document RAG tasks, limiting their applicability to scenarios requiring the processing of complex document corpora. To bridge this gap, we introduce DocRAGLib, a comprehensive dataset specifically tailored for multi-document RAG tasks.
The documents of DocRAGLib are derived from two primary sources: existing public single-document QA datasets and web-scraped heterogeneous data. First, we incorporate samples from MultiHiertt (zhao-etal-2022-multihiertt), where each document comprises textual content and hierarchical tables (averaging 3.89 tables per document). Second, we scrape contextual information utilizing webpage links provided in HiTab (cheng-etal-2022-hitab), reconstructing heterogeneous documents that contain both text and tables. This hybrid collection strategy ensures broad domain diversity, covering multiple fields such as finance, social, and healthcare, as shown in Figure 2.
In total, DocRAGLib contains 2,178 documents, combining heterogeneous text and table content, and includes 4,468 QA pairs meticulously designed based on these documents. Further details regarding Data Cleaning and additional statistical analysis are elaborated in Appendix A.
Figure 2. Distribution of domains in DocRAGLib
4. Methodology
@@
Figure 3. The Overview of our MixRAG Framework
@@
4.1. Heterogeneous Document Representation Module
The Heterogeneous Document Representation Module focuses on generating effective representations of heterogeneous documents, capturing both lengthy text and complex hierarchical tables to optimize retrieval performance. A traditional approach, the Table Level Summary, involves providing the table’s schema to the LLM and summarizing it. However, it oversimplifies the table, leading to the loss of key information and ignoring important structural details. For effective representation, the Heterogeneous Document Representation Module first introduces a row-and-column-level (RCL) table summary to represent tables. Then, leveraging the hierarchical structure of the tables, we extend this RCL table summary to capture the nested relationships and dependencies within the table data, ensuring a more comprehensive representation.
@@ 4.1. Heterogeneous Document Representation Module
(2) $\left\{\begin{aligned} H_{t}&=\{H_{t}^{r}\mid r=1,\dots,R_{t}\}\\ H_{l}&=\{H_{l}^{r}\mid r=1,\dots,R_{l}\}\end{aligned}\right$.
where $R_{t}$ denotes the number of hierarchical levels in the top headers, and $R_{l}$ denotes the number of hierarchical levels in the left headers.
@@
where $h_{t}^{r}(j)$ represents the header at level $r$ in the top header hierarchy, while $h_{l}^{r}(i)$ represents the header at level $r$ in the left header hierarchy. As shown in Figure 4, these headers form multi-level paths used to locate specific data cells within the table. To pinpoint a particular data cell $d_{ij}$, we define the paths from both the left and top headers, $P_{l}(i)$ and $P_{t}(j)$, respectively, which together uniquely identify the location of the data cell:
(4) $\left\{\begin{aligned} P_{l}(i)&=h_{l}^{1}(i_{1})\to h_{l}^{2}(i_{2})\to\dots\to h_{l}^{R_{l}}(i_{R_{l}})\\ P_{t}(j)&=h_{t}^{1}(j_{1})\to h_{t}^{2}(j_{2})\to\dots\to h_{t}^{R_{t}}(j_{R_{t}})\end{aligned}\right$.
@@
4.1.1. General RCL Table Summary